# Distillation of Large Language Models via Concrete Score Matching

### v1

@@ 4.2 Task-Specific Distillation
Experimental setup: We verify the effectiveness of CSD across diverse tasks, including dialogue summarization, low-resource translation, and arithmetic reasoning. Experiments were conducted on the DialogSum (Chen et al., 2021), Flores-200 (Costa-Jussà et al., 2022), and GSM8K (Cobbe et al., 2021) datasets. We used GEMMA-7B-IT (Team et al., 2024), fine-tuned with SFT as the teacher and GEMMA-2B-IT as the student, following the experimental setup of Xu et al. (2025). We compared with the baselines using the same teacher, changing only the loss function. Please see Appendix˜C for further details.
Results: Table˜3 presents a performance comparison across three tasks against the baseline loss functions. The proposed CSD objective function achieved the best performance on all three tasks under identical experimental conditions. In the arithmetic reasoning task, we observed multiple cases where certain losses resulted in zero accuracy. A case study in Tables˜7, 8, 9 and 10 in Appendix˜D shows that these models often generate excessively long reasoning steps without reaching a final answer, reflecting a failure to learn proper formatting; moreover, much of the reasoning itself was incorrect. As illustrated in Figure˜3(a), the losses TV, SKL, and SRKL exhibit mode-seeking tendencies, which we conjecture may have caused collapses into suboptimal modes depending on the task and model. In contrast, CSD obtained stable performance by applying ($T,S$) weighting in this case.
@@ 4.3 Ablation Studies
Table˜4 presents ablation studies on the weighting function choices in CSD and compares them with direct logit distillation (DLD). DLD accommodates only a single weighting function. When comparing DLD with $T$, $U$, and $S$ against CSD with ($T,T$), ($U,U$), and ($S,S$), respectively, our method consistently achieved higher average scores. As shown in Theorem˜2, we hypothesize that the broader solution space positively contributed to this improvement. Figure˜8 in Appendix˜D shows that DLD restricts solutions to those with a residual constant of zero, CSD adapts residual constants per token, providing evidence that it explores a broader solution set. Beyond performance, CSD provides a more flexible loss design space through two weighting functions. As illustrated in Figure˜3(a), replacing ($S,S$) with ($U,S$) or ($T,S$) reduces ROUGE-L but increases diversity, highlighting that CSD can adapt to tasks requiring either mode-covering or mode-seeking properties. Figure˜5 compares the method to resolve $\mathcal{O}(|{\mathcal{V}}|^{2})$ computational cost: 1) using analytic gradient computation from Theorem˜3 and 2) Monte Carlo sampling. Employing the analytic form led to faster training and improved convergence, demonstrating that the value of Theorem˜3 translates into empirical performance.
Table 4: Ablation on the logit-level loss design space. $T$, $U$, and $S$ denote teacher, uniform, and detached student probabilities. ROUGE-L scores are averaged over five seeds; best scores are in bold.
Loss $w_{1}(\cdot)$ $w_{2}(\cdot)$ Dolly Eval Self-Instruct Vicuna Eval Super-NI UnNI Avg. ($\uparrow$) $T$ - 0.09 0.07 0.18 0.07 0.06 0.09 DLD $U$ - 11.25 5.55 9.10 9.02 8.24 8.63 $S$ - 24.22 12.01 15.42 25.44 24.88 20.39 $T$ $T$ 6.82 4.24 9.16 4.53 4.83 5.91 $U$ $U$ 17.21 8.08 14.27 13.19 14.07 13.37 CSD $S$ $S$ 24.94 12.06 15.78 24.60 25.88 20.65 (Ours) $U$ $S$ 24.15 12.25 15.25 22.55 25.19 19.88 $T$ $S$ 22.77 10.62 14.06 18.81 21.71 17.59
Figure 5: Ablation between analytic gradient calculation (Eq.˜9) and Monte Carlo sampling for the $(S,S)$ weighting $\mathcal{L}_{\text{CSD}}$ calculation.
@@ Appendix B Related Works
The choice of discrepancy metric between teacher and student probability distributions is central to knowledge distillation for large language models (LLMs). Prior work has predominantly employed either forward KL divergence (Hinton et al., 2015) or reverse KL divergence (Gu et al., 2024). These divergences, however, exhibit distinct biases: forward KL is inherently mode-covering, while reverse KL is mode-seeking. Consequently, optimization under either measure imposes an unavoidable trade-off between fidelity and diversity. To address this limitation, recent studies have explored alternative measures, including (generalized) Jensen–Shannon divergence (Wen et al., 2023; Agarwal et al., 2024), adaptive KL divergence (Wu et al., 2025), and $\alpha$$\beta$ divergence (Wang et al., 2025). Complementarily, Ko et al. (2024) introduced skew KL and skew reverse KL divergences to improve optimization stability. Beyond the KL family, total variation distance has also been investigated (Wen et al., 2023). Broadly, existing approaches extend in two directions: (i) instantiating different generating functions within the $f$-divergence family, or (ii) constructing hybrid objectives that combine multiple divergences. In contrast, we propose a novel logit-level distillation framework grounded in concrete-score matching (Meng et al., 2022), which departs from the $f$-divergence family and offers both extensibility and originality. Furthermore, we introduce a loss design space with multiple instances, including instances that envelope the diversity–fidelity trade-off exhibited by previous methods.
Concurrently, a complementary line of work has examined dataset composition to mitigate the distribution mismatch between training and inference. Several studies have explored on-policy strategies, either using only student-generated outputs (Lin et al., 2020) or combining them with a fixed dataset (Agarwal et al., 2024) and teacher-generated outputs (Gu et al., 2024; Xu et al., 2025). To reduce the computational overhead of on-policy training, Ko et al. (2024) proposed an adaptive off-policy method with a replay buffer. By contrast, our contribution focuses on developing a novel discrepancy metric, which is orthogonal to these dataset composition strategies and can be seamlessly integrated with them as shown in Table˜2.
@@ Appendix C Experimental Details
Task-agnostic instruction-following distillation: All experiments were conducted primarily on four RTX 3090 GPUs. Following prior work (Gu et al., 2024; Ko et al., 2024; Wang et al., 2025), we searched learning rates in [5e-4, 1e-4, 5e-5] and batch sizes in [8, 16, 32]. Each configuration was trained for 20 epochs, saving a checkpoint at every epoch, and evaluated using the checkpoint with the highest validation ROUGE-L score. We used the same five evaluation seeds [10, 20, 30, 40, 50] as in prior work to compute the mean and standard deviation of the evaluation metric. The baselines in Table˜2 were run with the official code settings of prior work (Ko et al., 2024), with additional tuning for the batch size. In the OpenLLaMA experiments, all baselines and ours were standardized to a batch size of 8, the maximum supported in our environment. Baselines used the learning rates from their official code, while we fixed the learning rate to 1e-4 (commonly effective for GPT-2) with CSD, without further tuning. For ablation studies in Table˜4 and Figure˜5, we used the same configuration: learning rate 1e-4 and batch size 8. For GPT-4 feedback in Figure˜4, we use the following templates following prior work (Zheng et al., 2023; Ko et al., 2024) as shown below. We computed the ratio between the model answer and the golden answer for each of the 500 samples from Dolly Eval, and reported the average over all samples. We provide the reference implementation for CSD in Code LABEL:lst:CSD.
@@
Task-specific distillation: We follow the experimental setup of Xu et al. (2025) with a fixed dataset. For teacher SFT, we trained summarization and arithmetic reasoning for 3 epochs and translation for 10 epochs, using the full datasets. Model evaluation was performed every 16 steps, and the checkpoint with the lowest validation loss was selected.The batch size was fixed to 128 for all tasks, with the learning rate set to 1e-5. For each task in distillation, we distilled both the baselines and our method from the same teacher checkpoint with a fixed learning rate of 1e-5 and batch size of 8, using about 1,000 samples. We trained for 3 epochs on summarization and arithmetic reasoning, and 10 epochs on translation. For the baselines, checkpoints were saved every 25 steps, and the one with the lowest validation loss was used for evaluation. For CSD, since the loss itself cannot be directly computed and training relies on its gradient, validation loss was unavailable; thus, we evaluated using the final checkpoint. For all tasks, we set $w_{1}$ and $w_{2}$ using the teacher’s and student’s probabilities. For evaluation, we used task-specific metrics: COMET (Rei et al., 2022) for translation, ROUGE-L (Lin, 2004) for summarization, and answer accuracy for arithmetic reasoning, all evaluated on each task’s test dataset.

### v2

@@ 4.2 Task-Specific Distillation
We evaluate the effectiveness of CSD across dialogue summarization, low-resource translation, and arithmetic reasoning tasks. Distillation was conducted on 1,000 samples from the DialogSum (Chen et al., 2021), Flores-200 (Costa-Jussà et al., 2022), and GSM8K (Cobbe et al., 2021) datasets, following the experimental setup of Xu et al. (2025) (Please refer to Section˜C.2 for further details). Table˜3 compares performance across the three tasks against baseline loss functions. Under identical experimental conditions, the proposed CSD objective achieved the best results on all tasks. For the arithmetic reasoning task, we observed several cases in which certain losses yielded zero accuracy. A case study in Tables˜11, 12, 13, 14 and 15 of Appendix˜D shows that these models often produce excessively long chains of thought without arriving at a final answer, indicating a failure to learn proper formatting; furthermore, much of the reasoning itself is incorrect. As illustrated in Figure˜3(a), the RKL, TV, SKL, and SRKL losses exhibit mode-seeking tendencies, which we conjecture may have caused collapses into suboptimal modes under these limited data distillation settings. Similarly, CSD $(S,S)$, which also shows mode-seeking behavior, performed poorly as 21.09, 63.78, and 0.00 on summarization, translation, and GSM8K, respectively. CSD achieved stable performance in this case by using the $(T,S)$ weighting. DLD likewise performed poorly under student weighting and became only marginally stable when teacher weighting was applied. A full ablation is provided in Table˜10 of Appendix˜D for DLD variants. Across all cases, DLD remained significantly weaker than CSD, likely due to its restricted solution space being more detrimental under limited-data distillation.
Table 4: Benchmarking result for general chat capability. The best score is highlighted in bold.
 Qwen2.5-7B-IT → Qwen2.5-1.5B-IT  Gemma2-9B-IT → Gemma2-2B-IT
Benchmark  MT-Bench (0-10)  AlpacaEval (WR)  MT-Bench (0-10)  AlpacaEval (WR)
Judge  GPT4  GPT4-Turbo  GPT4-Turbo  GPT4  GPT4-Turbo  GPT4-Turbo
Teacher 8.59 7.52 88.69 8.91 7.66 94.60
DPKD (Li et al., 2024) 1.04 1.09 0.32 6.30 4.89 71.18
DistiLLM-2 (Ko et al., 2025) 7.28 5.75 70.42 7.81 6.45 89.91
DLD ($T$) 7.25 5.56 69.80 5.85 4.45 31.24
DLD ($S$) 7.28 5.74 67.67 7.58 6.53 89.84
CSD ($T,S$) 7.42 5.90 70.42 7.85 6.55 89.92
CSD ($S,S$) 7.69 5.95 69.64 7.77 6.43 90.05
4.3 General Chat Capability Distillation
To evaluate general chat performance, we conducted distillation experiments using the latest instruction-tuned models, Qwen2.5-Instruct (Team and others, 2024) and Gemma2-Instruct (Team et al., 2024). We followed DistiLLM-2 (Ko et al., 2025) and performed distillation using the 50K subset of the UltraChat dataset (Ding et al., 2023) (Please refer to Section˜C.3 for further details). Table˜4 shows that CSD outperformed both DistiLLM-2 and DLD on MT-Bench (Zheng et al., 2023) and AlpacaEval (Li et al., 2023) win rate (against text-davinci-003), demonstrating superior performance. These results further demonstrate the scalability of CSD.
@@ 4.4 Ablation Studies
CSD vs DLD: Table˜5 presents comprehensive ablation studies on the weighting function choices in CSD and compares them with direct logit distillation. DLD accommodates only a single weighting function. When comparing DLD with $T$, $U$, and $S$ against CSD with ($T,T$), ($U,U$), and ($S,S$), respectively, our method consistently achieved higher average scores. As shown in Theorem˜2, we hypothesize that the broader solution space positively contributed to this improvement. We also compared CSD against other shift-aware DLD variants such as DLD-min and DLD-max, as well as ranking-matching variants including DLD-std (Sun et al., 2024) and DLD-range. However, none of these methods outperformed the naive DLD baseline. Figure˜5(a) shows the effect of temperature scaling on the weighting function of DLD (s). Under the same temperature scaling, CSD consistently outperformed DLD. Figure˜9 in Appendix˜D shows that DLD restricts solutions to those with a residual constant of zero; CSD adapts residual constants per token, providing evidence that it explores a broader solution set.
Design choice of CSD: CSD provides a more flexible loss design space through two weighting functions. While $(S,S)$ provides high-fidelity generation, ($U,S$) and ($T,S$) have its own benefits. As illustrated in Figure˜3(a), replacing ($S,S$) with ($U,S$) or ($T,S$) reduces ROUGE-L but increases diversity, highlighting that CSD can adapt to tasks requiring either mode-covering or mode-seeking properties. As shown in Figure˜13 in Appendix˜D, the ($U,S$) weighting substantially reduces the gradient concentration on a small subset of vocabulary tokens that is induced by softmax. This allows all vocabulary items to be learned more evenly, which in turn produces the pattern in Figure˜9(c) where the logits residuals are tightly centered around their offsets. When minority-vocabulary logits are well learned, performance improves noticeably in situations where their contribution becomes more significant, such as under high-temperature sampling. This explains why ($U,S$) performs exceptionally well in the high-diversity region of Figure˜3(b). Figure˜5(b) measures probability calibration performance. While $(S,S)$ provides high-fidelity generation, it becomes overconfident. Therefore, in scenarios where probability calibration is preferred, we recommend using the $(T,S)$ weighting. Better calibration can potentially improve training stability in small-data settings where the optimization landscape is more difficult, as in Section˜4.2. Figure˜5(c) compares the method to resolve $\mathcal{O}(|{\mathcal{V}}|^{2})$ computational cost: 1) using analytic gradient computation from Theorem˜3 and 2) Monte Carlo sampling. In both cases, the performance is far superior to KL, but the analytic CSD shows slightly faster training and better convergence. Thus, we recommend using the analytic gradient whenever possible. However, the Monte Carlo variant allows extensions such as joint weighting functions or using Lp loss instead of L2, making it a viable option for more complex modeling.
Table 5: Ablation on the logit-level loss design space using GPT-2-0.1B student. $T$, $U$, and $S$ denote teacher, uniform, and detached student probabilities. ROUGE-L scores are averaged over five seeds; best scores are in bold. Please refer to Section˜A.4 for DLD variants details.
[…]
(a) Weight scaling
(b) Calibration error
(c) Training curve
Figure 5: Ablation studies for logit-level distillation loss design space.
@@ 5 Conclusion
Acknowledgments
This work was supported by the IITP (Institute of Information & Communications Technology Planning & Evaluation)-ITRC (Information Technology Research Center) grant funded by the Korea government (Ministry of Science and ICT) (IITP-2026-RS-2024-00437268). (50 $\%$) This research was supported by AI Technology Development for Commonsense Extraction, Reasoning, and Inference from Heterogeneous Data(IITP) funded by the Ministry of Science and ICT(RS-2022-II220077) (50 $\%$)
@@ A.3 Proof of Theorem˜3
A.4 The variants of direct logit distillation
We define the DLD variants used for our comparisons in Table˜2, Table˜3, and Table˜5.
$\mathcal{L}^{\text{min}}_{\text{DLD}}\left(\theta;p_{T},w\right)=\frac{1}{2}\sum_{y_{t}\in\mathcal{V}}w(y_{t})\left((f_{\theta}[y_{t}]-\min_{x}f_{\theta}[x])-(f_{T}[y_{t}]-\min_{x}f_{T}[x])\right)^{2},$
$\mathcal{L}^{\text{max}}_{\text{DLD}}\left(\theta;p_{T},w\right)=\frac{1}{2}\sum_{y_{t}\in\mathcal{V}}w(y_{t})\left((f_{\theta}[y_{t}]-\max_{x}f_{\theta}[x])-(f_{T}[y_{t}]-\max_{x}f_{T}[x])\right)^{2},$
$\mathcal{L}^{\text{std}}_{\text{DLD}}(\theta;p_{T},w)=\frac{1}{2}\sum_{y_{t}\in\mathcal{V}}w(y_{t})\left(\frac{f_{\theta}[y_{t}]-\frac{1}{|\mathcal{V}|}\sum\limits_{x\in\mathcal{V}}f_{\theta}[x]}{\sqrt{\frac{1}{|\mathcal{V}|}\sum\limits_{x\in\mathcal{V}}\left(f_{\theta}[x]-\frac{1}{|\mathcal{V}|}\sum\limits_{x^{\prime}\in\mathcal{V}}f_{\theta}[x^{\prime}]\right)^{2}}}-\frac{f_{T}[y_{t}]-\frac{1}{|\mathcal{V}|}\sum\limits_{x\in\mathcal{V}}f_{T}[x]}{\sqrt{\frac{1}{|\mathcal{V}|}\sum\limits_{x\in\mathcal{V}}\left(f_{T}[x]-\frac{1}{|\mathcal{V}|}\sum\limits_{x^{\prime}\in\mathcal{V}}f_{T}[x^{\prime}]\right)^{2}}}\right)^{2},$
$\mathcal{L}^{\text{range}}_{\text{DLD}}\left(\theta;p_{T},w\right)=\frac{1}{2}\sum_{y_{t}\in\mathcal{V}}w(y_{t})\left(\left(\frac{2\left(f_{\theta}[y_{t}]-\min_{x}f_{\theta}[x])\right)}{\max_{x}f_{\theta}[x]-\min_{x}f_{\theta}[x]}-1\right)-\left(\frac{2\left(f_{T}[y_{t}]-\min_{x}f_{T}[x])\right)}{\max_{x}f_{T}[x]-\min_{x}f_{T}[x]}-1\right)\right)^{2},$
$\mathcal{L}^{\text{mean}}_{\text{DLD}}\left(\theta;p_{T},w\right)=\frac{1}{2}\sum_{y_{t}\in\mathcal{V}}w(y_{t})\left(\left(f_{\theta}[y_{t}]-\frac{1}{|V|}\sum_{x\in\mathcal{V}}f_{\theta}[x]\right)-\left(f_{T}[y_{t}]-\frac{1}{|V|}\sum_{x\in\mathcal{V}}f_{T}[x]\right)\right)^{2},$
$\mathcal{L}^{\text{w-mean}}_{\text{DLD}}\left(\theta;p_{T},w\right)=\frac{1}{2}\sum_{y_{t}\in\mathcal{V}}w(y_{t})\left(\left(f_{\theta}[y_{t}]-\sum_{x\in\mathcal{V}}w(x)f_{\theta}[x]\right)-\left(f_{T}[y_{t}]-\sum_{x\in\mathcal{V}}w(x)f_{T}[x]\right)\right)^{2}.$
Here, $\mathcal{L}^{\text{w-mean}}_{\text{DLD}}$ is recovered by CSD as a special case shown by the following remark.
Remark 4.
Let CSD objective as $\mathcal{L}_{\text{CSD}}\left(\theta;p_{T},w\right)$ with $w(y_{t},x)=w_{1}(y_{t})w_{2}(x)$. If $w_{1}(\cdot)=w_{2}(\cdot)$, $\nabla_{\theta}\mathcal{L}_{\text{CSD}}\left(\theta;p_{T},w\right)=\nabla_{\theta}\mathcal{L}^{\text{w-mean}}_{\text{DLD}}\left(\theta;p_{T},w_{1}\right)$.
Proof.
We use $\sum_{y_{t}\in\mathcal{V}}w_{1}(y_{t})\tilde{f}^{w_{1}}_{\theta}[y_{t}]=\sum_{y_{t}\in\mathcal{V}}w_{1}(y_{t})\left(f_{\theta}[y_{t}]-\sum_{x\in\mathcal{V}}w_{1}(x)f_{\theta}[x]\right)=0$.
$\nabla_{\theta}\mathcal{L}^{\text{w-mean}}_{\text{DLD}}\left(\theta;p_{T},w_{1}\right)=\sum_{y_{t}\in\mathcal{V}}w_{1}(y_{t})\left(\tilde{f}^{w_{1}}_{\theta}[y_{t}]-\tilde{f}^{w_{1}}_{T}[y_{t}]\right)\left(\nabla_{\theta}f_{\theta}[y_{t}]-\sum_{x\in\mathcal{V}}w_{1}(x)\nabla_{\theta}f_{\theta}[x]\right)$
$=\sum_{y_{t}\in\mathcal{V}}w_{1}(y_{t})\left(\tilde{f}^{w_{1}}_{\theta}[y_{t}]-\tilde{f}^{w_{1}}_{T}[y_{t}]\right)\nabla_{\theta}f_{\theta}[y_{t}]-\cancelto{0}{\left(\sum_{y_{t}\in\mathcal{V}}w_{1}(y_{t})\left(\tilde{f}^{w_{1}}_{\theta}[y_{t}]-\tilde{f}^{w_{1}}_{T}[y_{t}]\right)\right)}\left(\sum_{x\in\mathcal{V}}w_{1}(x)\nabla_{\theta}f_{\theta}[x]\right)$
$=\nabla_{\theta}\mathcal{L}_{\text{CSD}}\left(\theta;p_{T},w\right)$
A.5 Weighting for divergence-based loss
Unlike the L2 loss, a divergence-based loss does not guarantee convergence to the target distribution when a weighting is applied. Here, we examine the effect of applying $w$ weighting to the KL divergence. Define $p^{w}_{T}(y_{t})=\frac{p_{T}(y_{t})\times w(y_{t})}{Z_{wT}},$ where $Z_{wT}=\sum_{y_{t}\in{\mathcal{V}}}(p_{T}(y_{t})\times w(y_{t}))$ is a partition function. Then we have:
$D^{w}_{\text{KL}}\left(p_{T}||q_{\theta}\right)$ $:=\sum_{y_{t}\in\mathcal{V}}w(y_{t})p_{T}(y_{t})\log{\frac{p_{T}(y_{t})}{q_{\theta}(y_{t})}}.$
$=\sum_{y_{t}\in\mathcal{V}}w(y_{t})p_{T}(y_{t})\left(\log{\frac{w(y_{t})p_{T}(y_{t})}{q_{\theta}(y_{t})}}-\log{w(y_{t})}\right).$
$=Z_{wT}\times\sum_{y_{t}\in\mathcal{V}}\frac{w(y_{t})p_{T}(y_{t})}{Z_{wT}}\left(\log{\frac{w(y_{t})p_{T}(y_{t})}{Z_{wT}\times q_{\theta}(y_{t})}+\log{Z_{wT}}}-\log{w(y_{t})}\right)$
$=Z_{wT}\times\sum_{y_{t}\in\mathcal{V}}p^{w}_{T}(y_{t})\left(\log{\frac{p^{w}_{T}(y_{t})}{q_{\theta}(y_{t})}+\log{Z_{wT}}}-\log{w(y_{t})}\right)$
$=Z_{wT}\times\sum_{y_{t}\in\mathcal{V}}p^{w}_{T}(y_{t})\log{\frac{p^{w}_{T}(y_{t})}{q_{\theta}(y_{t})}}+C$
$=Z_{wT}D_{\text{KL}}\left(p^{w}_{T}||q_{\theta}\right)+C,$
where $C$ is constant with respect to $\theta$. Thus, in this case, the student $q_{\theta}$ does not converge to the teacher distribution but instead converges to $p^{w}_{T}$, meaning that the target distribution is altered by the weighting $w$. In contrast, the proposed CSD theoretically guarantees convergence to the target distribution for any choice of $w$ proved by Proposition˜1.
@@ Appendix B Related Works
Distillation loss: The choice of discrepancy metric between teacher and student probability distributions is central to knowledge distillation for large language models (LLMs). Prior work has predominantly employed either forward KL divergence (Hinton et al., 2015) or reverse KL divergence (Gu et al., 2024). These divergences, however, exhibit distinct biases: forward KL is inherently mode-covering, while reverse KL is mode-seeking. Consequently, optimization under either measure imposes an unavoidable trade-off between fidelity and diversity. To address this limitation, recent studies have explored alternative measures, including (generalized) Jensen–Shannon divergence (Wen et al., 2023; Agarwal et al., 2024), adaptive KL divergence (Wu et al., 2025), and $\alpha$$\beta$ divergence (Wang et al., 2025). Complementarily, Ko et al. (2024) introduced skew KL and skew reverse KL divergences to improve optimization stability. Beyond the KL family, total variation distance has also been investigated (Wen et al., 2023). Amid (Shin et al., 2026) extends the distillation objective by matching the student distribution to a generalized assistant distribution within the f-divergence framework. Broadly, existing approaches extend in two directions: (i) instantiating different generating functions within the $f$-divergence family, or (ii) constructing hybrid objectives that combine multiple divergences. In contrast, we propose a novel logit-level distillation framework grounded in concrete-score matching (Meng et al., 2022), which departs from the $f$-divergence family and offers both extensibility and originality. Furthermore, we introduce a loss design space with multiple instances, including instances that envelope the diversity–fidelity trade-off exhibited by previous methods.
Distillation dataset: Concurrently, a complementary line of work has examined dataset composition to mitigate the distribution mismatch between training and inference. Several studies have explored on-policy strategies, either using only student-generated outputs (Lin et al., 2020) or combining them with a fixed dataset (Agarwal et al., 2024) and teacher-generated outputs (Gu et al., 2024; Xu et al., 2025). To reduce the computational overhead of on-policy training, Ko et al. (2024) proposed an adaptive off-policy method with a replay buffer. By contrast, our contribution focuses on developing a novel discrepancy metric, which is orthogonal to these dataset composition strategies and can be seamlessly integrated with them as shown in Table˜2.
Autoregressive model meets score function: BPO (Kim et al., 2025b) generalizes the preference optimization objective for LLMs by interpreting it as sequence-level concrete score matching. Unlike CSD, which formulates concrete score matching at the per-token level over the entire vocabulary to match a student model to a teacher model, BPO defines concrete scores at the sequence level between preferred and dispreferred responses and matches them to preference data. To enable tractable learning from such data, BPO further extends the objective within a Bregman divergence framework. ARD (Kim et al., 2025a), in contrast, extends the distillation of score-based diffusion models to an autoregressive model. Specifically, it employs an autoregressive student to learn, through regression, the integration of the ODE defined by the score function.
@@ Appendix C Experimental Details
C.1 Task-agnostic Instruction Following Distillation in Section˜4.1.
Experimental setup: We follow the training setup of DistiLLM (Ko et al., 2024). For the distillation dataset $\mathcal{D}$, we use databricks-dolly-15k (Conover et al., 2023), containing about 14,000 samples for training, with 500 held out for validation and 500 for evaluation. For comparison with the baseline, we optionally add a pretraining loss using the pretraining dataset OpenWebText (Gokaslan and Cohen, 2019) in some cases of Table˜2. We first fine-tune the GPT-2-1.5B (Radford et al., 2019) teacher on the dataset, and then distill it into GPT-2-0.1B and GPT-2-0.3B students. Similarly, we distill OpenLLaMA-7B (Geng and Liu, 2023) into OpenLLaMA-3B. We determined the learning rate and batch size by referring to the search ranges used in prior studies (Gu et al., 2024; Ko et al., 2024). We use the detached student probability as the default choice for both $w_{1}$ and $w_{2}$, and analyze alternative choices through ablation studies.
@@
All experiments were conducted primarily on four RTX 3090 GPUs. We searched learning rates in [5e-4, 1e-4, 5e-5] and batch sizes in [8, 16, 32]. Each configuration was trained for 20 epochs, saving a checkpoint at every epoch, and evaluated using the checkpoint with the highest validation ROUGE-L score. We used the same five evaluation seeds [10, 20, 30, 40, 50] as in prior work to compute the mean and standard deviation of the evaluation metric. The baselines in Table˜2 were run with the official code settings of prior work (Ko et al., 2024), with additional tuning for the batch size. In the OpenLLaMA experiments, all baselines and ours were standardized to a batch size of 8, the maximum supported in our environment. Baselines used the learning rates from their official code, while we fixed the learning rate to 1e-4 (commonly effective for GPT-2) with CSD, without further tuning. For ablation studies in Table˜5 and Figure˜5, we used the same configuration: learning rate 1e-4 and batch size 8. For GPT-4 feedback in Figure˜4, we use the following templates following prior work (Zheng et al., 2023; Ko et al., 2024) as shown below. We computed the ratio between the model answer and the golden answer for each of the 500 samples from Dolly Eval, and reported the average over all samples. We provide the reference implementation for CSD in Code LABEL:lst:CSD.
C.2 Task-specific Distillation in Section˜4.2.
Experimental setup: We verify the effectiveness of CSD across diverse tasks, including dialogue summarization, low-resource translation, and arithmetic reasoning. Distillation was conducted on DialogSum (Chen et al., 2021), Flores-200 (Costa-Jussà et al., 2022), and GSM8K (Cobbe et al., 2021) datasets, following the experimental setup of Xu et al. (2025) with a fixed dataset. We used Gemma-7B-IT (Team et al., 2024), fine-tuned with SFT as the teacher and Gemma-2B-IT as the student. We compared with the baselines using the same teacher, changing only the loss function.
@@
For teacher SFT, we trained summarization and arithmetic reasoning for 3 epochs and translation for 10 epochs, using the full datasets. Model evaluation was performed every 16 steps, and the checkpoint with the lowest validation loss was selected. The batch size was fixed to 128 for all tasks, with the learning rate set to 1e-5. For each task in distillation, we distilled both the baselines and our method from the same teacher checkpoint with a fixed learning rate of 1e-5 and batch size of 8, using about 1,000 samples. We trained for 3 epochs on summarization and arithmetic reasoning, and 10 epochs on translation. For the baselines, checkpoints were saved every 25 steps, and the one with the lowest validation loss was used for evaluation. For CSD, since the loss itself cannot be directly computed and training relies on its gradient, validation loss was unavailable; thus, we evaluated using the final checkpoint. For all tasks, we set $w_{1}$ and $w_{2}$ using the teacher’s and student’s probabilities. For evaluation, we used task-specific metrics: COMET (Rei et al., 2022) for translation, ROUGE-L (Lin, 2004) for summarization, and answer accuracy for arithmetic reasoning, all evaluated on each task’s test dataset.
C.3 General Chat Capability Distillation in Section˜4.3.
Experimental setup: We closely followed the official code of DistiLLM-2 (Ko et al., 2025) for our distillation setup. For 50k UltraChat prompts, we generated samples from both the student and the teacher. We then applied the CSD and DLD matching losses to each pair of generated samples. Although methods such as DPKD (Li et al., 2024) and DistiLLM-2 define different losses depending on which model produced the sample, CSD could also benefit from adopting such a strategy, suggesting room for further improvement. For evaluation, we followed the SimPO (Meng et al., 2024) protocol. We used the same learning rate and batch size as DistiLLM-2.
@@
Input: Student $f_{\theta}$, teacher $f_{T}$, prompt $\mathbf{c}$}, prefix ${\mathbf{y}}_{<t}$, function $w(\cdot,\cdot)=w_{1}(\cdot)w_{2}(\cdot|\cdot)$.
1 Compute the student logit $f_{\theta}[y_{t}]=f_{\theta}({\mathbf{c}},{\mathbf{y}}_{<t})[y_{t}],\forall y_{t}\in\mathcal{V}$.
2 Compute the teacher logit $f_{T}[y_{t}]=f_{T}({\mathbf{c}},{\mathbf{y}}_{<t})[y_{t}],\forall y_{t}\in\mathcal{V}$.
3 Sample $y_{t}$ according to $w_{1}(\cdot)$.
4 $\mathcal{L}^{\text{MC}}_{\text{CSD}}\left(\theta;p_{T},w\right)=\sum_{x\in\mathcal{V}}[w_{2}(x|y_{t})\times(f_{\theta}[x]-f_{\theta}[y_{t}]-f_{T}[x]+f_{T}[y_{t}])^{2}]$
return $\nabla_{\theta}\mathcal{L}^{\text{MC}}_{\text{CSD}}\left(\theta;p_{T},w\right)$
Algorithm 2 Monte Carlo estimation to compute $\mathcal{L}_{\text{CSD}}$ in Eq.˜8